Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2015-10-12

Observers: Heather Campbell, Joe Lyman
TAT: Janet Chen, Annalisa De Cia
Data reducers: Morgan Fraser

Lost the first half of the night to clouds. Second half remained thin/clear.
Wind 5-8m/s, seeing variable and reaching above 1.5" for parts.


UT      Target                  Comment
========================================================================
23:30                           Cloudy conditions, dome closed.
04:40                           Clouds cleared, dome opened + image analysis
04:50   ASASSN-15lh             Followup (gr16 2x1800s)
06:00   PS15cem                 Followup photometry (UBVRi)
06:20   PSNJ0109...             Followup (gr13 1500s, 1.5" slit)
07:00   ASASSN-15qc             Followup (g11+16 @900s, 1.5" slit)
07:40   ASASSN-15og             Followup photometry (BVRi)
07:55   L745a                   Spec standard
08:25   ASASSN-15qz             Classification (II; not in marshall yet)
08:40   PS15cel                 Classification (II)
09:00   PS15cgh                 Classification (pushed deep into twilight, noisy but shows blue continuum)
09:50                           Sky flats BR


Comments